Please help me calculate percent yield and theoretical yield for this experiment.

*0,3031g 1,4-dimethoxybenzene,

*0.5 mL of t-butanol

-T butanolDensity: 781 kg/m³

* 1.0 mL glacial acetic acid

-Density acetic acid: 1.05 g/cm³

*1.0 mL concentrated sulfuric acid

-Density Sulfuric acid:1.83 g/cm³

*15 mL of ice water

*product before crystallization: 0,3000g

*Product crystallized: 0,0812 g

Thanks
Answer #1

Since two equivalent of tert-butanol is used in the reaction therefore two tert-butyl are introduced in the aromatic moiety.

Molecular weight of reactant = 138.07

Molecular weight of product = 250.19

Therefore theoretical yield of reaction = (0.3031*250.19)/138.07

Theoretical yield of reaction = 0.549 gm

Product obtained before crystallization = 0.30 gm

Product obtained after crystallization = 0.0812 gm

Practical yield before crystallization = (0.3/0.549)*100 = 54.64%

Practical yield after crystallization = (0.0812/0.549)*100 = 14.79%